OpenStudy (anonymous):

evaluate if a=2 and b=-5 (8b)^2

OpenStudy (anonymous):

i dont get it

OpenStudy (angela210793):

sorry i misunderstood it =8*(-5)=-40 (-40)^2=1600

OpenStudy (anonymous):

thanks i got it

OpenStudy (anonymous):

\[1600\]